### v1

# Examining Algorithmic Curation on Social Media: An Empirical Audit of Reddit’s r/popular Feed

Abstract
Platforms are increasingly relying on algorithms to curate the content within users’ social media feeds. However, the growing prominence of proprietary, algorithmically curated feeds has concealed what factors influence the presentation of content on social media feeds and how that presentation affects user behavior. This lack of transparency can be detrimental to users, from reducing users’ agency over their content consumption to the propagation of misinformation and toxic content. To uncover details about how these feeds operate and influence user behavior, we conduct an empirical audit of Reddit’s algorithmically curated trending feed called r/popular. Using 10K r/popular posts collected by taking snapshots of the feed over 11 months, we find that the total number of comments and recent activity (commenting and voting) helped posts remain on r/popular longer and climb the feed. Using over 1.5M snapshots, we examine how differing ranks on r/popular correlated with engagement. More specifically, we find that posts below rank 80 showed a sharp decline in activity compared to posts above, and that posts at the top of r/popular had a higher proportion of undesired comments than those lower down. Our findings highlight that the order in which content is ranked can influence the levels and types of user engagement within algorithmically curated feeds. This relationship between algorithmic rank and engagement highlights the extent to which algorithms employed by social media platforms essentially determine which content is prioritized and which is not. We conclude by discussing how content creators, consumers, and moderators on social media platforms can benefit from empirical audits aimed at improving transparency in algorithmically curated feeds.
1 Introduction
Social media platforms are flooded with immense amounts of new content every day. This constant stream of content has led to the reliance on algorithms to curate the content in users’ social media feeds. Algorithmic curation is defined as the process of “organizing, selecting, and presenting subsets of a corpus of information for consumption” (Rader and Gray 2015). In recent times, the most prominent examples of algorithmic curation are on short-form video platforms like TikTok, Instagram Reels, and YouTube Shorts. However, less thought of are the “hot” and “popular” feeds that exist on multiple platforms (e.g., GitHub, Reddit, X/Twitter) that also use algorithms to curate what is “trending.”
Despite the prominent usage of algorithmic curation on social media, uncovering what factors influence curation algorithms is challenging. This is because curation algorithms are proprietary, preventing those on the outside from knowing how these algorithms work, thus protecting companies’ intellectual property. Additionally, algorithmic curation on social media is mostly done for specific users, i.e., what these algorithms recommend changes for each individual user. The prominence of personalized recommendations on social media presents two challenges: (1) understanding how these systems affect users more broadly, and (2) collecting enough data to support studies into these systems. To alleviate these challenges, we present a study that examines the trending feed on Reddit called r/popular.
Reddit’s r/popular Feed
The r/popular feed111https://www.reddit.com/r/popular on Reddit is one of the default feeds that is available to all users, with and without an account. The posts on the r/popular feed come from nearly all consenting communities on the platform—however, there are some subreddits, particularly “not safe for work” (NSFW) communities, that are omitted. The r/popular feed is a ranked list (see Figure 1) where the rankings are based on a calculated “hot” score (likely based on recency, comment rate, and upvote score) as well as some measures for post and discussion quality.222www.reddit.com/r/changelog/comments/9n3ix9/rpopular˙is˙changing/ At a high level, the r/popular feed consolidates the most active posts on Reddit from numerous subreddits and serves as the front page to the platform.
Studying Reddit’s r/popular feed comes with several benefits: (1) it is highly visited, (2) it is available to all users, and (3) it orders the posts consistently for all users, making it impactful for broad population. Its prominence also allows us to collect significant amounts of data to uncover what goes into Reddit’s ranking decisions. Thus, analyzing the r/popular feed can provide key insights into how algorithmically curated feeds work and how algorithmic ranking decisions can impact user behaviors on platforms like Reddit.
Our Contributions
In this paper, we conduct an algorithmic audit of r/popular with two key objectives. The first is to understand what factors influence algorithmic ranking on r/popular. The second is to quantify how those decisions, specifically the feed’s ranking decisions, affect the engagement on posts. Toward these goals, we ask three research questions.
What factors affect how long a post stays on the r/popular feed (i.e., the post’s tenure on the feed)?
What factors affect the assigned rank/position of the post on the r/popular feed?
How does the post’s rank on the r/popular feed affect the engagement on the post?
RQ1 and RQ2 examine the factors that influence Reddit’s ranking algorithm on r/popular. Specifically, what factors affect how long a post stays and where it is placed on the feed. RQ3 aims to quantify how algorithmic ranking decisions affect subsequent engagement on posts.
To answer these research questions, we capture a snapshot of the r/popular feed every 2 minutes over an 11-month period. Using over 1.5M consistently collected snapshots, we employ multiple regression analyses to examine the activity and movements (i.e., changes in position on the r/popular feed) of 10K posts from 694 distinct subreddits.
Summary of Findings
Through our analyses, we find that the total number of comments, along with recent commenting and voting activity, were the most predictive factors for a longer stay on the r/popular feed (RQ1) and for upward movement on the feed (RQ2). We also find that undesired comments, i.e., toxic and moderator-removed comments, were also predictive of a longer stay and upward movement, but to a lesser degree. Regarding how ranking on r/popular affects engagement (RQ3), we find that posts which were higher on the feed received comments at a higher rate, as well as a greater proportion of undesired comments.
By systematically analyzing snapshots of r/popular, we provide insights into how a prominent, algorithmically curated trending feed made its ranking decisions and how those decisions may influence user engagement on the platform. Understanding how these opaque algorithms operate as an external party is inherently difficult. The internal workings of curation algorithms are proprietary and therefore inaccessible to users and researchers. This lack of access limits transparency regarding how user engagement is driven by content ranking and how the content users interact with is influenced by engagement. Our approach can empower users and researchers to examine the algorithmically curated feeds that determine what they interact with on social media.
2 Related Work
In this section, we review prior work on trending feeds/popularity, algorithmic curation systems, and algorithmic audits.
Popularity & Trending Feeds
Popularity on social media has been studied before; however, these studies often focus on how users (DeVito 2022), communities (Maldeniya et al. 2020; Chan et al. 2024), or topics (Schlessinger et al. 2023) change after becoming viral. For example, Gurjar et al. (2022) found that after users became viral, they increased their posting frequency and changed their posts to be similar to the post that made them viral. In general, the focus of these studies is the impact of popularity on (groups of) users. However, missing from prior work are investigations into why content or users became viral in the first place. To address this gap, we examine the factors that influence the systems that determine what is viral on social media, specifically the r/popular feed on Reddit. It is important to understand how these algorithmic curation systems work because they are pervasive on social media and the popularity they induce can be a double-edged sword, particularly for those of marginalized identities (DeVito 2022). Additionally, these systems can also propagate inflammatory content as noted by mainstream media in recent times (O’Sullivan 2019; Katherine J. Wu 2019; Faddoul, Chaslot, and Farid 2020).
Algorithmic Curation & Ranking
According to Rader and Gray (2015), algorithmic curation is the process of “organizing, selecting, and presenting subsets of a corpus of information for consumptions.” Within this definition, we are particularly focused on the organizing and presenting aspect because Reddit algorithmically ranks posts on r/popular. However, there exist other definitions (Cotter, Cho, and Rader 2017; Eckles 2022; He et al. 2023). Algorithmic ranking is only one key mechanism in which algorithmic curation systems can exert power (Diakopoulos 2014). By ranking content, specifically on social media feeds, these algorithms essentially have the power to determine what is important by prioritizing content or users over one another. To show how impactful algorithmic ranking can be, Joachims et al. (2017) examined users’ clickthrough behavior on Google’s result page and found that participants’ trust in Google’s retrieval/ranking function led them to click on highly ranked links regardless of their quality or relevance to the query. Additionally, Salganik, Dodds, and Watts (2006) found that ranking songs in music markets by total downloads produced more unpredictability and inequality compared to groups who had songs ranked randomly. We extend this line of work by examining the influence of algorithmic ranking on social media trending feeds.
Algorithmic Audits
Per Metaxa et al. (2021), an algorithmic audit is “a method of repeatedly and systematically querying an algorithm with inputs and observing the corresponding outputs in order to draw inferences about its opaque inner workings.” In this paper, we query the r/popular feed for 11 months and examine its ranking decisions to infer details about its internals and impact on engagement. Our study complements the large corpus of existing social media algorithm audits performed on platforms like X/Twitter (Wang et al. 2024), YouTube (Ribeiro et al. 2020; Liu, Wu, and Resnick 2024), Facebook (González-Bailón et al. 2023), and TikTok (Mousavi, Gummadi, and Zannettou 2024) to name a few. Additionally, prior studies often focus on political bias that may be built into curation algorithms which differs from our focus on the factors that influence, and are influenced by, popularity. We extend this line of work and conduct an audit of algorithmic ranking on a relatively underexplored site, Reddit’s r/popular feed.
To conduct these algorithmic audits, researchers commonly utilize sock-puppet accounts (Bartley et al. 2021; Bandy and Diakopoulos 2021; Mousavi, Gummadi, and Zannettou 2024) that “use code scripts to create simulated users” (Liu, Wu, and Resnick 2024). Sock-puppet accounts are necessary because platforms often do user-specific recommendations based on the user’s activity. However, that raises the challenge of simulating realistic user behavior which is a limitation to these types of studies. By studying a trending feed that is available to everyone, we avoid having to use sock-puppet accounts.
3 Data
To retrieve posts from the r/popular feed, we used PRAW—a Python wrapper for Reddit’s API. Extending Chan et al. (2024)’s methodology, every 2 minutes, we captured a snapshot of the r/popular feed’s top 100 posts from March 23, 2022, to February 8, 2023—approximately 11 months. A single API request returns at most 100 posts. Although the feed goes beyond the top 100, we settled on only requesting the top 100 posts to avoid having to make multiple requests for a single snapshot and to stay within Reddit API rate limits. Thus, during the study period, we made only one request every 2 minutes.
Sampling Approach
By taking snapshots of r/popular every 2 minutes, totaling 224,121 feed snapshots, we collected 134,661 unique posts from 1,423 distinct subreddits. For clarity, feed snapshots refer to a capture of the entire feed whereas a snapshot from now on refers to a capture of an individual post within a feed snapshot of which there are 22,412,100 snapshots—100 per feed snapshot because there are 100 posts in a feed snapshot.
For tractability, we randomly sampled 10,000 posts from the 134,661 posts observed. The resulting sample contains 694 subreddits and 1,548,266 snapshots. For the rest of the paper, our findings are based on this representative sample of 10,000 posts and their respective snapshots. From the sample, we found that posts stay on r/popular for about 164 snapshots on average (, ). Additionally, posts from the sample, on average, stay on r/popular’s top 100 for 6.1 hours (, ). Along with these snapshots, we used Pushshift (Baumgartner et al. 2020) to obtain the comments for each post.
Identifying Undesirable Activity
In recent years, news outlets have suggested that social media platforms are intentionally promoting antisocial content through algorithmic prioritization to drive greater user engagement (Katherine J. Wu 2019; O’Sullivan 2019). To investigate whether antisocial behavior has any interactions with algorithmic curation, specifically on r/popular, we employed Almerekhi, Kwak, and Jansen (2022)’s fine-tuned BERT model, which assigns three toxicity-related scores for each comment: non_toxic, slightly_toxic, and highly_toxic. Each score ranges from 0 to 1. If a comment exceeded a score of 0.5—a threshold used in prior work (Bao et al. 2021)—for either slightly_toxic or highly_toxic, then we labeled it as toxic. Almerekhi, Kwak, and Jansen (2022) fine-tuned the model using r/AskReddit comments and showed its generalizability to 99 other large subreddits. Since the subreddits that appear on r/popular are mostly large, we claim that using this model is well suited for identifying toxic comments made within our r/popular posts. Additionally, some comments were removed, presumably by a moderator or bot, before they could be archived by Pushshift. To better capture antisocial behavior, we combine comments that contained “[removed]” with the ones flagged by the BERT model under an umbrella term: “undesired comments.”
Features
In this section, we describe the features we used for our regression models and provide a brief justification for inclusion. This feature list applies to posts captured at a specific snapshot/time. Thus, the features are measured at a specific time, e.g., the number of comments at a particular snapshot.
The first feature is:
1.
Content Type: Whether the post contains a link, video, image, or just text. This is a categorical variable where image posts are the reference category, i.e., the category in which all other content types are compared to.
Content type functions as a control variable to capture differences between posts that include images (49.19%), links (16.19%), text (12.45%), and videos (22.17%).
2.
Rank: Where the post is on r/popular where rank 1 is the top of the feed and rank 100 is the bottom.
3.
Age (hours): The amount of time, in hours, since the post’s creation.
Rank is the focus of our audit and because r/popular emphasizes content that is currently popular, the recency of a post is a natural feature to include.
The next set captures the activity within the post’s thread, i.e., comment section. We also have features that utilized the labels produced by the BERT model described previously.
4.
Num. Comments: The total number of comments at the time the snapshot was taken.
5.
Recent Comments: The number of comments made in the previous 10 minutes.
6.
Proportion Undesired: The proportion of comments that were labeled as undesired comments.
7.
Proportion Recent Undesired: The proportion of comments labeled as undesirable in the last 10 minutes.
8.
Score: The number of upvotes on the post minus the number of downvotes.
9.
Recent Votes: The number of votes the post received in the last 10 minutes.
10.
Proportion Upvotes: The proportion of all votes that were upvotes.
We included recent activity because there is a “hot” calculation2 that likely includes comment and vote velocity, i.e., how many comments are coming in currently. A 10-minute window was selected because posts stay at a rank for an average 7.5 of minutes before moving onto another (, ). Calculating the proportion of undesired comments helps us test whether Reddit’s ranking algorithm is in any way influenced by the presence of undesired activity.
11.
Num. Subscribers: The number of subscribers the post’s origin subreddit has.
Lastly, communities are integral to Reddit, which is why we included their size as a control variable in the models.
Table 1 provides the geometric mean and standard deviations for each feature. We used the geometric versions of these measures because the variables are log-transformed in our regression analyses. Additionally, the geometric standard deviations are used to inform the units we used to scale the regression coefficients which are also shown in Table 1. The consistency to use 2x for all units except the proportion of upvotes is to assist with interpretability.
| Feature | Unit | Median | ||
| Age (hours) | 2x | 8.738 | 1.644 | 9.067 |
| Num. Comments | 2x | 745.569 | 3.130 | 864 |
| Rec. Comments | 2x | 11.829 | 3.164 | 12 |
| Prop. Undesired | 2x | 0.189 | 1.660 | 0.190 |
| Rec. Prop. Undesired | 2x | 0.160 | 2.417 | 0.201 |
| Score | 2x | 13.161K | 2.587 | 13.697K |
| Rec. Votes | 2x | 357.792 | 3.021 | 396 |
| Prop. Upvotes | 1.05x | 0.917 | 1.065 | 0.930 |
| Num. Subscribers | 2x | 3.781M | 3.704 | 3.411M |
4 RQ1: Tenure on r/popular
To recall, RQ1 asks what factors are associated with how long a post stays on r/popular (i.e., tenure). To conduct this analysis, we built logistic regression models that estimate whether the post continues to exist on the feed in the next snapshot—approximately 2 minutes later. This section describes the model and its findings in further detail.
Model 1: Logistic Regression
Essentially, the task is to take an r/popular post during a snapshot and predict whether the same post continues to be on r/popular during the next snapshot. Thus, factors that helped a post stay on the feed in the next snapshot also helped elongate its tenure on r/popular. Because we did not observe posts outside the top 100, we built three models for the top 50 (), top 25 (), and top 10 () to have enough observations of posts both inside and outside of these intervals. Additionally, testing multiple rank intervals helped examine the robustness of our results. Each respective model only includes posts that, at some point, reached its respective rank interval (see Table 2 for the exact number of posts). For example, if a post is in snapshots before it reaches the top 50 and snapshots after, then uses only the snapshots after its initial breakthrough.
Results
| Feature | |||
| Link (vs. image) | 88.93%* | 9.77%* | -18.32%* |
| Text (vs. image) | 47.98%* | 11.12%* | -2.21% |
| Video (vs. image) | -0.13% | -6.24%* | 12.01%* |
| Age | -38.91%* | -39.16%* | -43.14%* |
| Num. Comments | 91.52%* | 105.37%* | 84.11%* |
| Rec. Comments | 19.08%* | 23.71%* | 27.76%* |
| Prop. Undesired | 2.64%* | 5.26%* | 16.80%* |
| Prop. Rec. Undesired | 3.18%* | 2.93%* | 3.94%* |
| Score | -47.89%* | -52.84%* | -57.45%* |
| Rec. Votes | 161.14%* | 125.10%* | 74.51%* |
| Prop. Upvotes | -0.60% | 4.10%* | -4.72%* |
| Num. Subscribers | 3.01%* | 2.39%* | 10.76%* |
| Num. Posts | 5,697 | 3,499 | 1,875 |
| Num. Snapshots | 1,062,717 | 687,386 | 391,167 |
| 0.230 | 0.261 | 0.285 |
Table 2 presents our findings from the three logistic regression models we built. The percentages in Table 2 correspond to the change in odds of staying in the top 50, 25, or 10 when the respective feature increases multiplicatively by its standardized unit listed in Table 1.
Impact of Overall Engagement. We found the strongest factors that increased the odds of staying near the top of the feed were the number of total comments, recent comments, and recent votes. We also found age (i.e., time since post creation) to be a strong factor that decreased the odds of staying near the top. These are consistent across all rank intervals and are natural for an engagement-based popular feed.
We also observed that increasing the score on a post corresponded to a drastic decrease in the odds of staying on the feed. We further discuss how this counterintuitive result may indicate the correlation between a post’s score and some unobserved influence in Section 8.
Impact of Undesired Activity. We found that increasing the proportion of undesired comments and the number of subscribers had a comparatively small but significant effect on increasing the odds that a post stays within the top 50 and top 25, with a stronger effect in the top 10. While these effects do not necessarily indicate that the ranking algorithm is intentionally designed to promote antisocial content to drive more user engagement, these results fail to rule it out. These effects also hint at an interaction between undesired comments and the rank of a post, and that the effects of undesired comments are strongest for posts at the top of the feed.
5 RQ2: Rank on r/popular
In this section, we describe our analyses of the factors associated with changes to a post’s rank on r/popular. As seen in Figure 2, a post does not gradually change between nearby ranks over time. Instead, the typical rank trajectory of a post consists of two parts: (1) where the post stays at one rank for several minutes, and (2) where the post suddenly jumps to another, distant rank. To characterize this discontinuous behavior, we used two separate regression models to determine which factors affect: (1) when a post changes rank, and (2) to which rank it changes when it does.
Model 2.1: Multinomial Regression
The first model is a multinomial regression that estimated whether a post will move up the feed, down the feed, or stay in its current position in the next snapshot. Posts tend to stay at the same rank for several minutes at a time (see Figure 2), and this model helped us determine which factors accelerated the movement of a post, and in which direction.
Model 2.2: Ordinal Regression
The second model is an ordinal regression that estimated the post’s next rank. Only snapshots where a post changed rank were considered so that we could get a more detailed picture of which rank a post changed to when it moved. We interpreted this model as a latent-variable model, where a post’s rank is determined by a continuous latent variable (), and the observed rank depends on which cut points falls between. The model estimated both the cut points and the association between each feature and the expected value of . The flexibility of the cut points within the model makes it well suited to handle the uneven “distances” between ranks—for e.g., it might be less likely for a post to move from rank 2 to rank 1 than it is from rank 99 to 98. We implemented this model in a Bayesian framework with strongly informative priors on the cut points and weakly informative priors on the coefficients. We provided more details on the priors in Appendix A.
| Feature | Up | Down |
| Link (vs. image) | 2.757% | 4.812%* |
| Text (vs. image) | 0.401% | 7.842%* |
| Video (vs. image) | 1.147% | -0.727% |
| Age | 4.117%* | 10.114%* |
| Num. Comments | -1.997%* | -7.200%* |
| Rec. Comments | 7.333%* | -4.127%* |
| Prop. Undesired | 0.751% | -2.681%* |
| Prop. Rec. Undesired | 0.285% | -0.801% |
| Score | -11.626%* | 11.018%* |
| Rec. Votes | 14.796%* | 3.886%* |
| Prop. Upvotes | 1.393%* | 0.803% |
| Num. Subscribers | 0.097% | -2.937%* |
| Num. Observations | 1,548,266 | |
| Nagelkerke | 0.043 |
Results
| Feature | |
| Link (vs. image) | 0.001 |
| Text (vs. image) | 0.033** |
| Video (vs. image) | -0.015* |
| Age | 0.163** |
| Num. Comments | -0.053** |
| Rec. Comments | -0.062** |
| Prop. Undesired | -0.021** |
| Prop. Rec. Undesired | -0.008** |
| Score | 0.132** |
| Rec. Votes | -0.081** |
| Prop. Upvotes | -0.018** |
| Num. Subscribers | -0.028** |
Starting with the multinomial regression model, Table 3 shows how a 2 times increase to a feature, excluding the proportion of upvotes, affected the probability of a post moving up or down the feed. Additionally, Figure 3 shows the baseline probabilities for each rank movement across the ranks on the feed. Note that the probability of moving up and down do not sum to 100% because the remaining amount corresponds to the probability that the post will not move in the next snapshot.
Impact of Rank on Movement. We found that posts closer to the top of the feed moved less frequently and less far. Figure 3 shows that the probability that a post moves in either direction reduces as its rank gets closer to 1, and Figure 4 that shows the gaps between cut points close to rank 1 are wider than cut points closer to rank 100. This is consistent with the intuition that there would be less competition toward the top of the feed as it would be rare for posts to achieve that level of “quality.”
Impact of Overall Activity on Rank. Similar to our results in RQ1, the number of recent comments and recent votes were strong factors that increased the probability that a post moved up in rank, as seen in Table 3. In terms of how far a post moved, Table 4 indicates that the total number of comments, recent comments, and recent votes had a similar magnitude effect. However, Table 3 indicates that the total number of comments tended to “stabilize” the rank of a post, decreasing the odds of movement in either direction, but still favoring upward movement. Similarly, the number of subscribers did not seem to increase the probability of upward movement directly, but it still favored upward movement by decreasing the probability of downward movement. These findings are consistent with an engagement-based ranking algorithm, but they highlight the influence of recent activity.
Impact of Undesired Activity on Rank. Table 3 indicates that the proportion of undesired comments slightly decreased the likelihood of moving down the feed, and Table 4 indicates that undesired comments had relatively little impact on where a post jumped in rank to. Thus, while undesired comments had little impact on moving posts up the feed, it reduced the rate at which posts move down.
For reasoning about score’s effect on rank, see Section 8.
6 RQ3: Engagment on r/popular
Our analyses from RQ1 and RQ2 provided empirical insights into the factors associated with ranking on r/popular. Finally, for RQ3, we investigated how rank on the r/popular feed influences the engagement on posts, specifically the frequency of comments gained before the next snapshot.
Model 3: Negative Binomial Regression
Our goal was to investigate the impact of features in Section 3 on commenting rate, and further, how they impacted the rate of undesired comments differently. We measured commenting rate as the number of comments gained between two consecutive snapshots. For example, if a post had 100 comments during a snapshot and at the next snapshot it had 120, then it has gained 20 comments. Since snapshots are taken approximately 2 minutes apart, this would correspond to a commenting rate of 10 comments per minute.
Since we wanted to investigate if there are differences in how features are associated with undesired commenting rate, we first consider the non-undesired commenting rate to be the reference which we compared the undesired commenting rate against. We used a single negative binomial regression model that estimated both non-undesired and undesired commenting rates using a dummy variable, which indicates that we are estimating undesired commenting rates. Thus, the model directly estimated the association between the features and the non-undesired commenting rate, and differences between how features are associated with undesired compared to non-undesired commenting rate are captured in interaction terms with the dummy variable. In other words, these interaction terms allowed us to inspect whether any of the features from Section 3 or the placement on r/popular has differing effects on undesired comments versus non-undesired comments.
Results
| Feature | Und. | Und:Und. | Net Und. |
| Link (vs. image) | -3.440%* | +0.342% | -3.110% |
| Self (vs. image) | -2.179%* | +0.650% | -1.543% |
| Video (vs. image) | 0.924%* | +0.006% | 0.930% |
| Age | -13.140%* | -0.467% | -13.545% |
| Num. Comments | 22.331%* | -2.601%* | 19.149% |
| Rec. Comments | 61.740%* | +1.056%* | 63.449% |
| Prop. Undesired | -14.467%* | +115.633%* | 84.437% |
| Prop. Rec. Und. | -1.500% | +8.713%* | 7.082% |
| Score | -13.992%* | +2.062%* | -12.218% |
| Rec. Votes | 13.530%* | -3.422%* | 9.645% |
| Prop. Upvotes | 3.519%* | -1.077%* | 2.405% |
| Num. Subscribers | 1.414%* | -0.344%* | 1.066% |
| Pseudo | 0.29 |
Table 5 shows how our feature set from Section 3, excluding rank which is visualized in Figure 5, is associated with the rate of non-undesired comments, the reference, the rate of and undesired comments.
Impact of Engagement on Non-Undesired Activity. From this table, we found that the number of comments, recent commenting activity, and recent voting activity had the strongest associations with future non-undesired activity. Additionally, the impact of recent commenting activity (61.74%) on non-undesired commenting rate was greater than the one found with the total number of comments (22.33%). This hints that the active discussions played a more important role to future engagement than the total number of discussions that have already occurred.
Impact of Engagement on Undesired Activity. For undesired activity, we found that the proportion of undesired comments had the strongest association (84.43%) out of all the ones found in Table 5. This suggests that having a greater proportion of undesired activity would invite similarly undesired activity in the future.
Impact of Rank on Commenting Activity. Regarding rank’s associations with commenting rate, we found that the rate of non-undesired comments dropped fairly consistently from ranks 2 through 80, visualized in the first subplot in Figure 5. Afterward, the rate of non-undesired comments fell precipitously below rank 80.
Impact of Rank on Undesired Activity. From the second subplot in Figure 5, we observed that the ratio of undesired comments fell as the post is placed lower on the feed. Specifically, the ratio of undesired comments to non-undesired comments fell at a rate of 0.064% per rank (). Additionally, ranks 2 and 3 may indicate that there is a sharp rise in the ratio at the peak of the r/popular feed, however, it is difficult to be sure given the confidence intervals. Overall, we found that posts higher on the r/popular feed had a slightly greater tendency to attract undesired comments compared to posts below it. It is also important to note that the direction of causation is unclear as we previously found that undesired comments may have a slight impact on rank as well.
7 Discussion
Here, we discuss the implications of our methodology and findings for future audits of algorithmically curated feeds.
Empowerment Through Transparency
Prior literature has shown that people are often unaware that algorithms control what they see and, in turn, how their content is spread (Rader, Cotter, and Cho 2018; Hsu et al. 2020). This unawareness can lead to harm (Eslami et al. 2015) resulting in calls for greater transparency regarding these algorithmic systems (Alvarado and Waern 2018). Our study enables transparency by quantifying how factors like commenting rate and undesirable activity affect algorithmic ranking on a prominent social media feed.
Implications for Content Creators. Our study has implications for content creators, particularly those whose livelihoods are tied to algorithmic decision making. The relationship we identified between algorithmic rank and engagement highlights how the algorithms employed by social media platforms essentially get to decide which content is prioritized and which is not. This has led content creators on platforms like YouTube to blindly attempt to appease algorithms by adjusting their content creation process (Choi et al. 2023). Audits of algorithmic feeds can empower creators by providing accurate information about how their actions can affect content prioritization on their respective platforms.
Implications for Content Consumption.
We found that the order in which content is ranked can influence the levels and types of user engagement within algorithmically curated feeds. Similarly, the types of content that are recommended and where they are positioned can affect users’ content consumption practices. Prior work has examined this impact through users’ trust in Google’s ranking algorithm, which leads to a greater number of clicks on top-ranked links regardless of their accuracy or relevance to the query (Joachims et al. 2017). This trust bias that Joachims et al. (2017) refer to, coupled with the increased engagement levels observed in our analysis (see Section 6), can catalyze the spread of misinformation—especially if higher ranks within feeds are misconstrued as an indicator of content quality and trustworthiness. To alleviate some of these pitfalls, some platforms have provided explanations for their algorithmic recommendations, however, Mousavi, Gummadi, and Zannettou (2024) found that those on TikTok in particular were generic and inapplicable to the user. Further research is needed to qualitatively understand how algorithmic rank affects users’ perceptions of content highlighted on social media feeds.
Implications for Content Moderation
Prior work has shown that sudden popularity can be disruptive to moderation teams that have to manage increased activity, particularly from newcomers (Kiene, Monroy-Hernández, and Hill 2016; Chan et al. 2024). In addition to increased activity levels, we found that posts ranked higher on r/popular corresponded to higher proportions of undesirable activity, which can make handling popularity even more challenging. Despite how integral algorithmic curation is to social media platforms, moderators currently have little influence on algorithmic curation systems. This lack of agency impedes their ability to direct their communities and highlight desirable content—an idea that has been explored in prior work (Choi et al. 2024). Given that moderators have little influence on these systems, there is a need for design interventions that provide moderators with more controls or ways to “override” algorithmic curation systems. These additional affordances can be used to highlight desirable content in feeds, which has theoretical foundations (Kraut and Resnick 2012) and is currently done with existing mechanisms such as awards, upvotes, pins, and flairs (Lambert, Choi, and Chandrasekharan 2024). These additional controls may take the form of various sliders that adjust latent weights on the feed to emphasize different priorities (e.g., increase user diversity on the feed), visual indicators for moderator-assigned contributions, or specific feeds that filter for moderator-selected contributions—similar to the ones the New York Times has to filter for reader- and editor-selected comments (Wang and Diakopoulos 2022).
Implications for Future Empirical Audits
Despite the pervasiveness of algorithmic curation on social media, studying it as external researchers is exceedingly difficult. This difficulty is compounded by data access restrictions as platforms lock down their APIs, e.g., X/Twitter (Calma 2023) and Reddit (Reddit 2023). The reduction in API accessibility has led researchers to use other methods like sock-puppet accounts (Perriam, Birkbak, and Freeman 2020; Bartley et al. 2021; Bandy and Diakopoulos 2021; Liu, Wu, and Resnick 2024; Mousavi, Gummadi, and Zannettou 2024) and calls for data donations (Jhaver et al. 2023; Chouaki et al. 2024)—each with their own set of drawbacks. These restrictions have led to the reliance on externally-maintained datasets like the now defunct Pushshift—which has been heavily used in prior research (Jhaver, Bruckman, and Gilbert 2019; Atcheson, Koshy, and Karahalios 2024; Kou et al. 2024; August et al. 2024). To avoid these challenges, we developed a robust pipeline to collect large-scale high-fidelity snapshots of Reddit’s trending feed r/popular which can be adapted for other platforms with similar feeds.
However, we recognize that these approaches are extremely time and resource-intensive, posing significant challenges for scalability without adequate financial backing and collaborative support. Although partnerships between academia and industry have the potential to yield important research outcomes, such collaborations are increasingly rare. Given these circumstances, and echoing a theme identified in a 2024 CCC Workshop Visioning Report (Eslami et al. 2024), this presents an opportunity for researchers to share auditing infrastructures and data-sharing protocols that support the collection, storage, and access of social media data for research purposes. For example, in the industry circuit, data clean rooms (Herbrich 2022) are emerging for several companies to have a shared and secure data infrastructure. However, the creation of these infrastructures must also adhere to relevant privacy and legal standards such as GDPR (GDRP 2023) or CCPA (CCPA 2018), especially for data that contain sensitive information that would need to be anonymized. This would enable more research like ours, informing future regulation, while safeguarding the rights and privacy of social media users.
8 Limitations
We recognize that our study bears limitations, however, these limitations also suggest interesting future directions.
Our findings are inherently tied to the demographics of Reddit users. Therefore, any attempt to apply our conclusions to other platforms should carefully consider the demographic similarities between Reddit and the platform in question.
The relatively low our models exhibited indicate that there are other factors outside of the ones we have included that influence the outcomes. One such factor may include details about surrounding posts on the r/popular feed. A hint that our model is incomplete is the observation that increasing the score on a post corresponds to a drastic decrease in the odds of staying on the feed, as found in RQ1 and RQ2. This counterintuitive result may be due to a correlation between a post’s score and some unobserved influence. Regardless, future work can build upon these models by including more factors about surrounding posts that are competing for the same finite number of spots on these feeds. Furthermore, future work could also employ more advanced statistical methods like a stochastic transitivity model (Johnson and Kuhn 2013) or a causal framework used in prior work (Chandrasekharan et al. 2017; Saha, Chandrasekharan, and De Choudhury 2019; Jhaver, Rathi, and Saha 2024). These causal approaches will help examine the causal relationship between the different factors that influence and are influenced by algorithmic ranking.
Lastly, because our findings were based on observational data, future work will have to examine more user-centered effects of algorithmic ranking—similar to Salganik, Dodds, and Watts (2006); Joachims et al. (2017), but on social media. Specifically, future work can structure controlled experiments that test the impacts of algorithmic rank on individual users and their perceptions of highly ranked content. These experiments can additionally include interviews or surveys to assess affective characteristics like trust, engagement, and perceived fairness. Together, these studies can offer complementary insights that provide a more comprehensive understanding of algorithmic ranking on social media platforms.
9 Conclusion
Now that algorithmic curation has become integral to online ecosystems, examining it and its effects on user behavior becomes critical. In this paper, we conducted a comprehensive empirical audit of one such system: Reddit’s r/popular feed. Through this, we successfully quantified how recent commenting rate and other factors influence algorithmic ranking on r/popular, as well as how rank/position on r/popular correlates with subsequent engagement and undesirable behavior. Our findings are based on millions of snapshots of the top 100 ranks on r/popular consistently collected over 11 months, an approach that can be applied to other platforms. Additionally, we discussed the implications of our findings for stakeholders, including content creators, highlighted moderators’ lack of agency in community curation systems, and proposed future research directions on algorithmic curation amid reduced data access. All in all, studying algorithmic curation goes back to user agency: should users have control over how they interact with content and how their content is distributed online? To address this, we must first understand how these systems are embedded in our online environments, thus enabling us to make informed decisions about governing them.
Appendix A Bayesian Priors for Model 2.2
Our intuition is that it is “easier” to move between ranks further down the feed (e.g., from 95 to 91) than it is to move between ranks higher up the feed (e.g., from 5 to 1). Indeed, based on Figure 6, we saw that the variance of rank in the next snapshot appears to increase approximately linearly with the rank in the current snapshot, up until rank 60-80. Beyond that, the variance appears to decrease due to the fact that ranks beyond 100 were censored from our dataset. We used this observation to set strongly informative priors on our cut points.
First, we assumed that the variance of next rank increases linearly with the current rank, then the standard deviation would increase proportionally to the square root of the current rank. In other words, this means that posts tend to traverse more ranks the further down the feed, and the number of ranks they typically traverse increases proportionally to the square root of the current rank. We can capture this notion in the priors for our cut points by making the gap between cut points inversely proportional to the square root of rank. Thus we chose the following priors for our cut points for each rank :
is a scaling factor with a weakly informative prior of , and is fixed to .
We also assume that a post’s next rank will be close to its current rank. Thus, the intercepts for each current rank were given weakly informative priors of .
Finally, the remaining coefficients were given weakly informative priors: